Requirement: Answers to each of the following problems will be evaluated based on accuracy, completeness and clarity. Unsupported answers will receive no credit. Any assumptions you make in answering the questions should be clearly stated. Point allocation is given beside each question.

Condition: Suppose Oregon proposes indexing the minimum wage to inflation. In the space below, summarize what it would mean to index the minimum wage to inflation, and then describe the substitution and scale effects you anticipate with this policy? (In your response, assume that the minimum wage is an effective price floor and that both factor and product markets are perfectly competitive.)

Question: Describe the anticipated substitution effect in this scenario.
Answer #1

Question 1) Answer :- False.

Explanation :- Profit- maximization condition for firm is Marginal revenue (MR) = Marginal cost (MC).

Q = 140 - P (Demand function)

P = 140 - Q (Inverse demand function)

Marginal revenue (MR) = 140 - 2Q ( The slope of marginal revenue function is twice that of the slope of inverse demand function.)

Marginal cost (MC) = 40

Equating marginal revenue (MR) and marginal cost (MC),

140 - 2Q = 40

2Q = 100

Q = 50

P = 140 - 50 = $ 90. (Put tha value of Q = 50 in inverse demand function.)

Total revenue = 50 * 90 = $ 4500

Total cost = 25 + 40 * 50 = $ 2025 (Put the value of Q = 50 in cost function given in the question.)

Accordingly, Profits = 4500 - 2025 = $ 2475.

The leader in stackelberg leadership model will earn profit of $ 2475 in the long run. Thus, the given statement in question that the leader under stackelberg leadership model will not make profits in long run is false. Even if followers play monopolistic competition, stackelberg leader will earn profits in long run.

Question 2). Answer :- False.

Explanation :- In stackelberg leadership model, leader firm move first and then the other follower firms move. The first move made by leader in stackelberg leadership model will help leader to make twice the profits made by follower firms. Accordingly, the given statement in question that leader and followers be in equal size is false because the market size of leader firm will be more than the follower firm due to more profits earned by leader firm as compared to other follower firms.